OpenStudy (anonymous):

Solve this system algebraically. 9x + 2y = 5 y - 2x + 3 = 0 {(11/13, -17/13)} {(-11/13, -17/13)} {(-11/13, 17/13)} {(11/13, 17/13)}

OpenStudy (shamim):

multiply ur second equation by 2 and then subtract

OpenStudy (austinl):

9x + 2y = 5 - 2x +y = -3

OpenStudy (austinl):

Multiply your second equation by negative 2. Add the top and bottom. 9x + 2y = 5 -2(- 2x +y = -3) 9x + 2y = 5 -4x - 2y = 6 5x = 11 x = 11/5 Then plug that back into one of your original equations.

OpenStudy (austinl):

x = 11/5 -2(11/5) + y = -3

OpenStudy (anonymous):

Ty

OpenStudy (austinl):

No problem!